OpenStudy (anonymous):

Factor the expression. x3 − 27

Parth (parthkohli):

Hint:\[\rm \large a^3 - b^3 = (a - b)(a^2 + ab + b^2)\]

OpenStudy (anonymous):

\[a^3-b^3 = (a-b) (a^2+ab+b^2)\]

OpenStudy (anonymous):

\[x^3 - 27 = x^3-3^3\]

OpenStudy (anonymous):

= (x)^3 - (3)^3 use formula that is: (a)^3 - (b)^3 = (a-b)(a^2 +ab + b^2) here you have a= x and b=3.. can you solve it now..??

OpenStudy (anonymous):

no i dont get it

OpenStudy (anonymous):

\[(x ^{2}+3x+9)(x-3)\]